1.    Given y = (x, z) = 6x2 + 3xz4 where

X(B, E, T) = B2 + 2E + TB

Z(B, E, T) = (ETB)1 + B + E Then, find the partial derivatives of y with respect to B, E, and T.


1
Expert's answer
2021-08-30T14:35:46-0400

Given that,

"y = (x, z) = 6x^2 + 3xz^4 \\where\\

X(B, E, T) = B^2 + 2E + TB\\

Z(B, E, T) = (ETB)1 + B + E \\

So,\\

y=6(B^2+2E+TB)^2+3(B^2+2E+TB)(ETB+B+E)^4"

Integrating partially with respect to B,

"\frac{∂y}{∂B}=\frac{∂}{∂B}[6(B^2+2E+TB)^2]+\frac{∂}{∂B}[6(B^2+2E+TB)(ETB+B+E)^4]\\=12(B^2+2E+TB)(2B+T)+3(2B+T)(ETB+B+E)^4+12(B^2+2E+TB)(ETB+B+E)^3(ET+1)"

Integrating partially with respect to E,

"\frac{∂y}{∂E}=\frac{∂}{∂E}[6(B^2+2E+TB)^2]+\frac{∂}{∂E}[6(B^2+2E+TB)(ETB+B+E)^4]\\=12(B^2+2E+TB)×2+3(ETB+B+E)^4×2+3(B^2+2E+TB)×4(ETB+B+E)^4(TB+1)\\=24(B^2+2E+TB)+6(ETB+B+E)^4+12(B^2+2E+TB)(ETB+B+E)^4(TB+1)"


Integrating partially with respect to T,

"\frac{∂y}{∂T}=12(B^2+2E+TB)B+3(ETB+B+E)^4B+12(B^2+2E+TB)(ETB+B+E)^3EB"